# Binary Stars Approaching Supermassive Black Holes: Tidal Break-up, Double Stellar Disruptions and Stellar Collision

### v1

@@ Abstract
In galactic centers, stars and binaries can be injected into low-angular-momentum orbits, resulting in close encounters with the central supermassive black hole (SMBH). We use $N$-body simulations to study such encounters systematically under a wide range of conditions. Depending on the system parameters (such as $\beta_{b}$, the ratio of binary tidal radius to pericenter distance $r_{p}$ to the SMBH, and the compactness of the binary), such close encounters can lead to the break-up of the binary, disruptions of both stars and collision between the stars. Binary break-up produces a hyper-velocity star and a bound star around the SMBH; the peak value of the orbital binding energy depends weakly on $\beta_{b}$. When $r_{p}$ is comparable to the stellar tidal radius, sequential disruptions of the stars occur within a time interval much shorter than the initial binary orbital period, potentially exhibiting distinct double TDE features. Stellar collisions occur for a range of $\beta_{b}$’s, with a few to 10’s percent probabilities (depending on the compactness of the binary). In gentle encounters ($\beta_{b}\lesssim 1$), stellar collisions occur after the pericenter passage, and the merger remnants are typically ejected from the SMBH at a small velocity. In deep encounters ($\beta_{b}\gtrsim 1$), collisions occur near the pericenter, and the merger remnants are typically bound to the SMBH. We suggest that stellar collisions induced by binary-SMBH encounters may produce exotic stars in galactic centers, trigger accretion flares onto the SMBH due to the mass loss, and result in bound merger remnants causing repeated partial TDEs.
@@ 1 Introduction
The unique conditions at the centers of galaxies, namely a large number of high-velocity stars in the deep gravitational potential well of a central supermassive black hole (SMBH), can lead to unusual patterns of star formation, dynamical evolution and nuclear activities (e.g., Alexander, 2017). Within the sphere of influence of the SMBH, two-body relaxation and other collective processes are efficient, and stars in the nuclear cluster can be “kicked” into low-angular-momentum orbits toward the SMBH, leading to tidal disruptions of the stars (Rees, 1988; Stone et al., 2020). Since the 1990s, many tidal disruption events (TDEs) have been detected (van Velzen et al., 2020; Gezari, 2021).
@@
Similar to TDEs, a stellar binary that passes sufficiently close to the SMBH is tidally disrupted, resulting in the ejection of a binary component with very high velocity (Hills, 1988). Many hyper-velocity stars (HVSs) flying away from the Galactic Center have been found (e.g. Brown et al., 2005). The velocities of these HVSs are so high (measured at $500-1000$ km/s) that binary break-ups by the Galactic Center SMBH provide a natural explanation (Yu & Tremaine, 2003; Gould & Quillen, 2003). A binary break-up also leaves behind a star on a bound, eccentric orbit around the SMBH. It has been suggested that such a star may be related to the phenomena of partial TDEs (Cufari et al., 2022; Somalwar et al., 2023) and quasi-periodic eruptions (Lu & Quataert, 2023; Linial & Sari, 2023; Zhou et al., 2024).
@@
Many previous works have studied the dynamics of close encounters between stellar binaries and SMBHs (e.g., Gualandris et al., 2005; Bromley et al., 2006; Ginsburg & Loeb, 2006; Sesana et al., 2007; Sari et al., 2009; Brown et al., 2018). In addition to the binary break-up that produces an HVS and a tightly bound star, other outcomes are possible. When the pericenter distance of the binary to the SMBH is comparable to the tidal radius of the star, both members of the binary will be disrupted sequentially; such double TDEs may exhibit light curves that are distinct from single TDEs (Mandel & Levin, 2015). During or shortly after the binary-SMBH close encounter, the two stars may come so close to each other that a physical collision occurs (Ginsburg & Loeb, 2007); such a collision leads to stellar merger and mass loss, which may produce exotic stars and generate observable electromagnetic signatures.
@@
This paper is devoted to a systematic study of the various possible outcomes of the close encounters between stellar binaries and SMBHs, including binary break-up, double stellar disruptions and stellar collision/merger. We extend previous works by surveying a wide parameter space, for encounter pericenter distances ranging from outside the binary’s tidal sphere to deep inside it. Using $N$-body simulations, we characterize the orbital energy distribution of the stars after the binary break-up, and determine the time separation between the two stellar disruptions in double TDEs. For stellar collisions, we are interested in questions such as (a) what is the probability of stellar collisions for different system parameters? (b) what is the velocity distribution of the collisions? (c) how likely are the merger remnants bound to the SMBH, and what are their orbital energies? Overall, an important goal of our study is to obtain the key results related to binary-SMBH encounters that can be scaled to different system parameters.
@@ 3 Binary Tidal Break-Up
The Hills mechanism for binary break-up is well understood (Hills, 1988). In this section, for completeness, we analyze the energy distribution of the stars produced by this mechanism.
@@
Figure 1: Probability density function (PDF) of the orbital energy of each star (in units of $E_{\rm tide}=E_{\rm tide1}$, given by Equation 14) after binary break-up. The blue and orange solid lines represent the results for two different values of $\beta_{b}$: gentle encounters ($\beta_{b}=0.7$) and deep encounters ($\beta_{b}=5$). These PDFs are obtained by sampling all encounters with different phase angles and inclinations. Note that the PDF does not normalized to unity, because for a given $\beta_{b}$, only a fraction of the encounters lead to binary break-ups; each PDF includes all encounters, not just those leading to binary break-ups.
@@ 4 Double stellar disruptions
Figure 2 shows our numerical results (assuming $m_{1}=m_{2}=m$) for the distribution of $\Delta t_{\rm DD}$ for encounters with $\alpha=8,\beta_{b}=20$ (Note that these parameters violate the condition in Equation 13). To test the mass scaling, we consider $M/m=10^{4},10^{5}$ and $10^{6}$. We see that there is indeed an upper bound on $\Delta t_{\rm DD}$ in the simulation data and that the upper bound agrees with Equation (15). The distribution of $\Delta t_{\rm DD}$ is nearly uniform between 00 and $(\Delta t_{\rm DD})_{\rm max}/2$ and decreases to zero towards $(\Delta t_{\rm DD})_{\rm max}$. Importantly, $\Delta t_{\rm DD}$ is much smaller than the orbital period of the binary (see Eq. 15). Since the occurrence of double stellar disruptions requires $\beta_{b}\gtrsim\alpha$, it tends to occur primarily for tight binaries (small $\alpha$) with shorter orbital periods. These suggest that double stellar disruption flares occur on a rather short time separation and may be distinguishable. Additionally, if the two stellar disruptions occur almost simultaneously, the interaction between the resulting debris is likely and may give rise to unique electromagnetic radiation signals.
@@ 5 Stellar collision
We now examine stellar collisions resulting from close encounters between binaries and SMBHs. Previous work has shown that collisions can happen during deep ($\beta_{b}\gg 1$) encounters (Ginsburg & Loeb, 2007). For a binary moving around the SMBH over many orbits, the Lidov-Kozai effect can induce very high eccentricity in the binary orbit and thus stellar collisions (Antonini et al., 2010). Here we focus on stellar collisions induced by a single (“first”) close encounter between the binary and the SMBH. In Section 5.1 we determine the parameter space where the collision occurs and the corresponding fraction of collisions; in Section 5.2 we analyze the collision velocities; in Section 5.3 we examine the properties of the merger remnants.
@@ 5.3.2 Capture Fraction of Merger Remnants
Figure 9: Same as Figs. 3 and 5, except that the figure shows the fraction of the stellar merger remnants (among the stellar collisions) that are captured by the SMBH. Note that for $\beta_{b}=5$ and $10$, we only consider $\alpha>2^{-2/3}\beta_{b}$ (see Eq. 13) to avoid stellar tidal disruption.
@@ 6.1 Summary of Key Results
We have carried out a systematic study on the close encounters between stellar binaries and SMBHs. The binary (with a total mass $m_{12}$, initial circular orbit and semi-major axis $a_{b}$) approaches the SMBH (mass $M$) on a nearly parabolic orbit, with pericenter distance $r_{p}$ to the SMBH. In addition to binary break-ups, such encounters can also lead to double stellar disruptions (double TDEs) and stellar collisions/mergers. The occurrences of these different outcomes depend on the system parameters through three dimensionless ratios: $\beta_{b}=r_{\rm tide}^{b}/r_{p}$, where $r_{\rm tide}^{b}=a_{b}(M/m_{12})^{1/3}$ is the tidal radius of the binary; $\beta_{\star}=r_{\rm tide}^{\star}/r_{p}$, where $r_{\rm tide}^{\star}$ is the tidal radius of the star; and $\alpha=a_{b}/R_{\rm col}$, where $R_{\rm col}$ is the separation of the two stars at collision. (For a binary with identical stars, $R_{\rm col}=2R_{\star}$ and $\beta_{b}=2^{2/3}\alpha\beta_{\star}$.) Using $N$-body simulations, we characterize these different outcomes as a function of the three dimensionless ratios. Our key findings are:
1. Binary break-up ($\beta_{b}\gtrsim 1$) produces a hyper-velocity star (HVS) and a tightly bound star around the SMBH – the well-known Hills mechanism. We determine the energy distribution of the HVS and the bound star for different $\beta_{b}$ (see Figure 1). This energy distribution peaks at the characteristic “tidal” energy (Eq. 14) and differs from the energy distribution of the tidal debris produced when a single star is disrupted. Deeper encounters (with $\beta_{b}\gtrsim$ a few) have the potential to produce HVSs with somewhat larger ejection velocities.
2. Sequential disruptions of the binary stars (double TDEs) occur when $\beta_{\star}\gtrsim 1$. We determine the distribution of the time interval between the two disruptions (Figure 2) and show that it is much shorter than the initial orbital period of the binary. This suggests that the debris from the two disrupted stars is likely to collide and may produce distinct electromagnetic signatures. Further hydrodynamic simulations will be needed to clarify these signatures.
3. Stellar collisions can happen for both gentle encounters ($\beta_{b}\lesssim 1$) and deep encounters ($\beta_{b}\gtrsim 1$). The occurrence fraction depends on the compactness of the binary (or the parameter $\alpha$), but can reach a few to 10’s percent (Figs. 3-5).
@@ 6.2.1 Bound Stars from Binary Break-ups
Bound stars from binary break-ups have been invoked to explain the so-called repeated partial TDEs. Several candidate events have been found/claimed recently, including ASASSN-14ko (period 115.2 days; Payne et al., 2021, 2022, 2023; Huang et al., 2023), eRASSt J045650.3–203750 ($299\rightarrow 193$ days; Liu et al., 2023, 2024), AT2018fyk ($\sim 1200$ days; Wevers et al., 2023), RX J133157.6–324319.7 ($\sim 10000$ days; Hampel et al., 2022; Malyali et al., 2023), AT 2020vdq ($\sim 870$ days; Somalwar et al., 2023) and AT 2022dbl ($\sim 710$ days; Lin et al., 2024). However, in many cases, this interpretation should be considered tentative. Since $\alpha$ is greater than a few for realistic binaries, most binary break-ups would not lead to immediate stellar disruption, unless $r_{p}\lesssim r_{\rm tide}^{\star}$, or $\alpha/\beta_{b}\lesssim 1$. Equation (28) gives the orbital period of the bound star immediately after the binary break-up. Further orbital decay due to gravitational radiation can reduce the period as well as the pericenter distance of the bound star: During this decay, the orbit evolves according to (Peters, 1964)
@@
So a reduction of the pericenter distance by a factor of a few would imply a reduction of the eccentricity to $\lesssim 0.2$. It is not clear whether a TDE with such a low eccentricity would be similar to the standard high-eccentricity TDEs.

### v2

@@ Abstract
In galactic centers, stars and binaries can be injected into low-angular-momentum orbits, resulting in close encounters with the central supermassive black hole (SMBH). Previous works have shown that under different conditions, such close encounters can lead to the break-up of the binary, disruptions of both stars and collision between the stars. We use 3-body scattering experiments to characterize these different outcomes for a range of system parameters, such as $\beta_{b}$, the ratio of binary tidal radius to pericenter distance $r_{p}$ to the SMBH and the compactness of the binary. We focus on stellar collisions, which occur for a range of $\beta_{b}$’s, with a few to 10’s percent probabilities (depending on the compactness of the binary). In gentle encounters ($\beta_{b}\lesssim 1$), stellar collisions occur after the pericenter passage, and the merger remnants are typically ejected from the SMBH at a small velocity. In deep encounters ($\beta_{b}\gtrsim 1$), collisions occur near the pericenter, with the impact velocity a few times the escape velocity of the star, and the merger remnants are typically bound to the SMBH. We suggest that stellar collisions induced by binary-SMBH encounters may produce exotic stars in galactic centers, trigger accretion flares onto the SMBH due to the mass loss, and result in bound merger remnants causing repeated partial TDEs.
@@ 1 Introduction
The unique conditions at the centers of galaxies, namely a large number of high-velocity stars in the deep gravitational potential well of a central supermassive black hole (SMBH), can lead to unusual patterns of star formation, dynamical evolution and nuclear activities (e.g., Alexander, 2017). Within the sphere of influence of the SMBH, two-body relaxation and other collective processes are efficient, and stars in the nuclear cluster can be “kicked” into low-angular-momentum orbits toward the SMBH, leading to tidal disruptions of the stars (Rees, 1988; Stone et al., 2020). Since the 1990s, many tidal disruption events (TDEs) have been detected (Komossa, 2015; van Velzen et al., 2020; Gezari, 2021). Similar to TDEs, a stellar binary that passes sufficiently close to the SMBH is tidally disrupted, resulting in the ejection of a binary component with very high velocity (Hills, 1988). Many hyper-velocity stars (HVSs) flying away from the Galactic Center have been found (e.g. Brown et al., 2005; Koposov et al., 2019). The velocities of these HVSs are so high (measured at $500-1000$ km/s) that binary break-ups by the Galactic Center SMBH provide a natural explanation (Yu & Tremaine, 2003; Gould & Quillen, 2003). A binary break-up also leaves behind a star on a bound, eccentric orbit around the SMBH. It has been suggested that such a star may be related to the phenomena of partial TDEs (Cufari et al., 2022; Somalwar et al., 2023) and quasi-periodic eruptions (Lu & Quataert, 2023; Linial & Sari, 2023; Zhou et al., 2024).
@@
Many previous works have studied the dynamics of close encounters between stellar binaries and SMBHs (e.g., Gualandris et al., 2005; Bromley et al., 2006; Ginsburg & Loeb, 2006; Sesana et al., 2007; Sari et al., 2009; Brown et al., 2018). In addition to the binary break-up that produces an HVS and a tightly bound star, other outcomes are possible. When the pericenter distance of the binary to the SMBH is comparable to the tidal radius of the star, both members of the binary will be disrupted sequentially; such double TDEs may exhibit light curves that are distinct from single TDEs (Mandel & Levin, 2015; Bradnick et al., 2017). During or shortly after the binary-SMBH close encounter, the two stars may come so close to each other that a physical collision occurs (Ginsburg & Loeb, 2007; Antonini et al., 2010; Bradnick et al., 2017); such a collision leads to stellar merger and mass loss, which may produce exotic stars and generate observable electromagnetic signatures.
@@
This paper is devoted to a systematic study of the various possible outcomes of the close encounters between stellar binaries and SMBHs, including binary break-up, double stellar disruptions and stellar collision/merger. We extend previous works by surveying a wide parameter space, for encounter pericenter distances ranging from outside the binary’s tidal sphere to deep inside it. Using restricted 3-body scattering experiments, we characterize the orbital energy distribution of the stars after the binary break-up, and determine the time separation between the two stellar disruptions in double TDEs. For stellar collisions, we are interested in questions such as (a) what is the probability of stellar collisions for different system parameters? (b) what is the velocity distribution of the collisions? (c) how likely are the merger remnants bound to the SMBH, and what are their orbital energies? Overall, an important goal of our study is to obtain the key results related to binary-SMBH encounters that can be scaled to different system parameters.
@@ 2.1 Key Parameters and Scalings
Note that in this paper we are considering the first close encounter between a SMBH and a relatively tight binary, with $\alpha\lesssim 15$. Thus, unless otherwise noted, we will assume that the initial binary orbit is circular. This is reasonable because tidal interaction between the binary components is expected to have circularized the orbit. Indeed, Meibom & Mathieu (2005) found that the characteristic circularization period of solar-type MS stars is $8-15$ days (depending on the stellar ages), which corresponds to $\alpha=11-16$.
@@ 2.2 Simulation Setup
While most of this paper deals with initially circular binaries, for completeness, we will also present some results for binaries with initially finite eccentricity, $e_{b}$. When considering elliptical binary orbits, the outcome of an encounter additionally depends on the longitude of the ascending node of the binary orbit ($\Omega$). In our simulations, we sample $\Omega$ with a flat prior between 0 and $2\pi$. Furthermore, because the phase variation of an elliptical orbit is not uniform, we no longer employ a flat prior when sampling the initial orbital phase ($\lambda$). Instead, we use the orbital phase values derived from a flat prior in the orbital time over the interval from 00 to $P_{b}$, the binary orbital period.
@@ 3 Binary Tidal Break-Up
The Hills mechanism for binary break-up is well understood (Hills, 1988). In this section, for completeness, we analyze the binary break-up fraction and the energy distribution of the stars produced by this mechanism.
Figure 1: The fraction of binary break-ups as a function of the dimensionless parameter $\beta_{b}\equiv r_{\rm tide}^{b}/r_{p}$ for several different values of the initial binary eccentricity.
Figure 1 shows the fraction of binary break-ups following an encounter between the binary and the SMBH. We see that binaries on (initially) elliptical orbits are more susceptible to tidal break-up. Moreover, even when the pericenter distance is smaller than the binary tidal radius ($\beta_{b}>1$), the break-up fraction does not reach unity (see also Sari et al., 2009).
@@
Figure 2: Probability density function (PDF) of the orbital energy of each star (in units of $E_{\rm tide}=E_{\rm tide1}$, given by Equation 14) after binary break-up. The initial binary has $e_{b}=0$. The blue and orange solid lines represent the results for two different values of $\beta_{b}$: gentle encounters ($\beta_{b}=0.7$) and deep encounters ($\beta_{b}=5$). These PDFs are obtained by sampling all encounters with different phase angles and inclinations. Note that the PDF does not normalized to unity, because for a given $\beta_{b}$, only a fraction of the encounters lead to binary break-ups; each PDF includes all encounters, not just those leading to binary break-ups.
@@
Figure 3: Same as Figure 2, except showing only the positive part of the energy (the negative part is perfectly symmetric with respect to the positive part) and for different initial binary eccentricities. The left panel represents gentle encounters ($\beta_{b}=0.7$), while the right panel represent deep encounters ($\beta_{b}=5$).
Figure 3 shows the distributions of $E_{1}$ for binaries with different initial eccentricities. We see that for small (but finite) eccentricities, the peak of the distribution slightly decreases (compared to the $e_{b}=0$ case), while the distribution itself becomes broader. This can be explained by changing $(2a_{b1})$ in Eq. (14) to the range between $[2a_{b1}(1-e_{b})]$ and $[2a_{b1}(1+e_{b})]$. For larger eccentricities, the peak of the energy distribution becomes significantly lower.
@@ 4 Double stellar disruptions
Figure 4 shows our numerical results (assuming $m_{1}=m_{2}=m$) for the distribution of $\Delta t_{\rm DD}$ for encounters with $\alpha=8,\beta_{b}=20$ (Note that these parameters violate the condition in Equation 13). To test the mass scaling, we consider $M/m=10^{4},10^{5}$ and $10^{6}$. We see that there is indeed an upper bound on $\Delta t_{\rm DD}$ in the simulation data and that the upper bound agrees with Equation (15). The distribution of $\Delta t_{\rm DD}$ is nearly uniform between 00 and $(\Delta t_{\rm DD})_{\rm max}/2$ and decreases to zero towards $(\Delta t_{\rm DD})_{\rm max}$. Importantly, $\Delta t_{\rm DD}$ is much smaller than the orbital period of the binary (see Eq. 15). Since the occurrence of double stellar disruptions requires $\beta_{b}\gtrsim\alpha$, it tends to occur primarily for tight binaries (small $\alpha$) with shorter orbital periods.
Note that although $\Delta t_{\rm DD}$ is small, if the stars have different masses/radii, their disrupted bound streams will return to the pericenter on different timescales (Mandel & Levin, 2015). Indeed, the return time of the most bound debris is
$t_{\rm min}\sim{\pi M\over m}\left(\frac{R_{\star}^{3}}{2GM}\right)^{1/2}.$ (16)
For $R_{\star}\propto m^{0.8}$, we have $t_{\rm min}\propto m^{0.2}$, which means that for binaries with unequal masses, we may see a double-peaked burst. In addition, the interaction between the two stellar streams may give rise to dissipation and electromagnetic radiation signals (Bonnerot & Rossi, 2019).
@@ 5 Stellar collision
We now examine stellar collisions resulting from close encounters between binaries and SMBHs. Previous work has shown that collisions can happen during deep ($\beta_{b}\gg 1$) encounters (Ginsburg & Loeb, 2007). For a binary moving around the SMBH over many orbits, the Lidov-Kozai effect can induce very high eccentricity in the binary orbit and thus stellar collisions (Antonini et al., 2010; Bradnick et al., 2017). Here we restrict to stellar collisions induced by a single (“first”) close encounter between the circular binary and the SMBH.
Due to the excitation of eccentricity during close encounters, it is likely that a binary will have a finite eccentricity when it approaches the black hole multiple times, increasing the probability of stellar collisions (Antonini et al., 2010; Bradnick et al., 2017).
In Section 5.1 we determine the parameter space where the collision occurs and the corresponding fraction of collisions; in Section 5.2 we analyze the collision velocities; in Section 5.3 we examine the properties of the merger remnants.
@@ 5.1.2 Deep Encounters ($\beta_{b}\gtrsim 1$)
5.1.3 Eccentric Binaries
For completeness, here we examine the impact of the initial binary eccentricity on the collision fraction. Note that to avoid stellar collision before the binary-SMBH encounter, we require $a_{b}(1-e_{b})>R_{\rm col}$, or $\alpha>1/(1-e_{b})$.
Figure 8: Same as Figs. 5 and 7, except for different initial binary eccentricities with two selected $\beta_{b}$.
Figure 8 illustrates the effect of different initial binary eccentricities on the collision fraction in both gentle encounters ($\beta_{b}=0.6$) and deep encounters ($\beta_{b}=5$). We see that for $e_{b}\lesssim 0.2$, the collision fraction is similar to the $e_{b}=0$ case. However, for higher eccentricities ($e_{b}\sim 0.8$), the collision fraction can be much larger. This suggests that if multiple encounters occur between the binary and the SMBH before tidal circularization can sufficiently reduce the eccentricity, the fraction of binary collisions will increase significantly (see Antonini et al., 2010; Bradnick et al., 2017).
@@ 5.3.2 Capture Fraction of Merger Remnants
Figure 12: Same as Figs. 5 and 7, except that the figure shows the fraction of the stellar merger remnants (among the stellar collisions) that are captured by the SMBH. Note that for $\beta_{b}=5$ and $10$, we only consider $\alpha>2^{-2/3}\beta_{b}$ (see Eq. 13) to avoid stellar tidal disruption.
@@ 6.1 Summary of Key Results
We have carried out a systematic study on the close encounters between stellar binaries and SMBHs. The binary (with a total mass $m_{12}$, initial circular orbit and semi-major axis $a_{b}$) approaches the SMBH (mass $M$) on a nearly parabolic orbit, with pericenter distance $r_{p}$ to the SMBH. In addition to binary break-ups, such encounters can also lead to double stellar disruptions (double TDEs) and stellar collisions/mergers. The occurrences of these different outcomes depend on the system parameters through three dimensionless ratios: $\beta_{b}=r_{\rm tide}^{b}/r_{p}$, where $r_{\rm tide}^{b}=a_{b}(M/m_{12})^{1/3}$ is the tidal radius of the binary; $\beta_{\star}=r_{\rm tide}^{\star}/r_{p}$, where $r_{\rm tide}^{\star}$ is the tidal radius of the star; and $\alpha=a_{b}/R_{\rm col}$, where $R_{\rm col}$ is the separation of the two stars at collision. (For a binary with identical stars, $R_{\rm col}=2R_{\star}$ and $\beta_{b}=2^{2/3}\alpha\beta_{\star}$.) Using 3-body scattering experiments, we characterize these different outcomes as a function of the three dimensionless ratios.
Our main findings concern the stellar collisions in both gentle encounters ($\beta_{b}\lesssim 1$) and deep encounters ($\beta_{b}\gtrsim 1$). The occurrence fraction of collisions depends on the compactness of the binary (or the parameter $\alpha$), but can reach a few to 10’s percent (Figs. 5-7).
@@
Additionally, for completeness, we present the break-up fraction and energy distribution of the stars following a binary breakup for different $\beta_{b}$’s and initial binary eccentricities (see Figs 1-3). Eccentric binaries are more susceptible to tidal break-up. The energy distribution peaks at the characteristic “tidal” energy (Eq. 14) for circular and slightly eccentric binaries, but peaks at lower value for highly eccentric binaries, and differs from the energy distribution of the tidal debris produced when a single star is disrupted. We also explore the double TDE scenario, where we show that the time interval between successive stellar disruptions is always small, making the difference of the return times of the two disrupted stellar streams more significant.
@@ 6.2.1 Bound Stars from Binary Break-ups
Bound stars from binary break-ups have been invoked to explain the so-called repeated partial TDEs. Several candidate events have been found/claimed recently, including ASASSN-14ko (period 115.2 days; Payne et al., 2021, 2022, 2023; Huang et al., 2023), eRASSt J045650.3–203750 ($299\rightarrow 193$ days; Liu et al., 2023, 2024), AT2018fyk ($\sim 1200$ days; Wevers et al., 2023), RX J133157.6–324319.7 ($\sim 10000$ days; Hampel et al., 2022; Malyali et al., 2023), AT 2020vdq ($\sim 870$ days; Somalwar et al., 2023) and AT 2022dbl ($\sim 710$ days; Lin et al., 2024). However, in many cases, this interpretation should be considered tentative. Since $\alpha$ is greater than a few for realistic binaries, most binary break-ups would not lead to immediate stellar disruption, unless $r_{p}\lesssim r_{\rm tide}^{\star}$, or $\alpha/\beta_{b}\lesssim 1$. Equation (29) gives the orbital period of the bound star immediately after the binary break-up. If the orbit is in the regime where the two-body scattering time is shorter than the gravitational radiation timescale, the eccentricity may continue to increase, potentially leading to a TDE. On the other hand, if the gravitational radiation timescale is shorter, the orbit will decay (Sari & Fragione, 2019), and evolve according to (Peters, 1964)
@@
So a reduction of the pericenter distance by a factor of a few would imply a reduction of the eccentricity to $\lesssim 0.2$. A TDE with such a low eccentricity would be different from the standard high-eccentricity TDEs.